# Galactic Archaeology with Gaia

### v1

@@ Abstract
The Gaia mission has revolutionized our view of the Milky Way and its satellite citizens. The field of Galactic Archaeology has been piecing together the formation and evolution of the Galaxy for decades, and we have made great strides, with often limited data, towards discovering and characterizing the subcomponents of the Galaxy and its building blocks. Now, the exquisite 6D phase-space plus chemical information from Gaia and its complementary spectroscopic surveys has handed us a plethora of data to pour over as we move towards a quantitative rather than qualitative view of the Galaxy and its progenitors. We review the state of the field in the post-Gaia era, and examine the key lessons that will dictate the future direction of Galactic halo research.
@@ 1 Introduction
Gaia is an astrometric mission with the lofty goal of producing an all-sky phase-space map of the Galaxy (Gaia Collaboration et al., 2016b). Positions on the sky, parallax, and proper motions are measured for all stars down to a limiting magnitude of $G\approx 21$. In addition, a (bright) subset of stars has measured radial velocities and astrophysical parameters (such as effective temperature and surface gravity). As of February 2024, we are now on the third Gaia data release with the fourth release likely arriving in 2025/26. Each subsequent data release provides data over a larger time baseline, and hence more precise astrometry, as well as additional data products, such as variability information and source classifications. To compensate for the bright limit for accurate Gaia spectroscopic measurements, several wide-field spectroscopic surveys have (or will) complemented the Gaia data to provide radial velocities and chemical measurements, e.g. SDSS/SEGUE (Yanny et al., 2009), LAMOST (Cui et al., 2012), APOGEE (Majewski et al., 2017), H3 (Conroy et al., 2019), DESI (Cooper et al., 2023), WEAVE (Jin et al., 2023), 4MOST (de Jong et al., 2019). These auxiliary datasets are particularly important for the halo stars, which traverse to large distances and are thus often relatively faint. The era of Gaia is truly data-heavy! The fruits of this labour will last for many years to come, and we are only at the beginning stages of our digestion and understanding of what this detailed mapping of the Milky Way contains.
The purpose of this review is to provide the current state of understanding of the Galactic halo and the revelations and surprises from the Gaia data boon. This review does not intend to provide a complete history and consensus, as the number of papers in this field has exploded in recent years. Instead, we focus on the main findings, in our humble opinion, and look ahead to the future of this field over the coming decade.
@@ 1.1 A dominant stellar halo progenitor?
One of the main aims of the Gaia mission is to piece together the building blocks of the Galactic halo. This ambitious goal was perhaps mainly expected to reveal a plethora of streams, clouds, and debris from a multitude of progenitors (see e.g. Helmi and de Zeeuw, 2000). However, one of the clearest and most decisive discoveries from the early Gaia data releases was that in fact the (inner) stellar halo is dominated by one dwarf galaxy progenitor (see Section 2). As we will discuss below, in hindsight this was not a surprising discovery!
@@ 1.1.1 Predictions from $\Lambda$CDM
The fundamental mass function of dark matter haloes ($dn/dM$ — i.e. how many haloes there are of different mass) follows a power-law profile $dn/dM\propto M^{-1.9}$ (e.g. Moore et al., 1999; Jenkins et al., 2001; Gao et al., 2004; Tinker et al., 2008). Thus, the general expectation is that there are many more smaller lumps than bigger lumps — a prediction that lies at the heart of the hierarchical $\Lambda$CDM paradigm. A consequence of the mass function is that haloes like our own Milky Way engulf many more smaller mass objects than more massive ones over time. For example, $N$-body simulations (Fakhouri et al., 2010) predict that a Milky Way-mass halo ($10^{12}M_{\odot}$) typically undergoes $N\sim 30$ minor mergers (mass-ratio between 1:3 and 1:100) and $N\sim 3$ major mergers (mass-ratio $>1:3$) since redshift $z=10$.
@@
When galaxies enter the equation, things become more complicated as the efficiency of star formation is not equal amongst haloes of different mass. The inefficiency of massive galaxies in the centres of clusters (mainly due to AGN feedback, e.g. Croton et al. 2006; Hopkins et al. 2006; Fabian 2012) and dwarf galaxies (supernova feedback plus reionization, e.g. Dekel and Silk 1986; Efstathiou 1992; Bullock et al. 2000; Pontzen and Governato 2012) at forming stars means that galaxies like the Milky Way lead the pack — we live in the most efficient star-making galaxy environment (see e.g. Figure 2 of Wechsler and Tinker, 2018). The relation between stellar mass and dark matter mass (aptly known as the ‘stellar mass-halo mass’ or SMHM relation), has important consequences for the growth of stellar haloes through the accretion of lower-mass objects. The SMHM has a very steep decline at low halo masses ($\lesssim 10^{11}M_{\odot}$, see e.g. Figure 6 of Bullock and Boylan-Kolchin 2017). Thus, although there may be hundreds of dark matter subhaloes accreted by the galaxy, if these have very few stars (or no stars at all!) then they have an insignificant contribution to the accreted stellar halo. This idea was recognised in Purcell et al. (2007), who used analytic prescriptions for dark matter halo growth and galaxy formation models, and later by Deason et al. (2016) using $N$-body simulations and analytic SMHM relations. In short, the contribution of stellar mass to the galactic stellar halo is dominated by a small number (1-2) of relatively massive dwarfs ($M_{\rm DM}\sim 10^{11}M_{\odot}$, $M_{\rm star}\sim 10^{8-9}M_{\odot}$). The contribution of lower mass dwarfs, while potentially important for dark matter growth, is relatively insignificant for the total stellar halo mass (see Fig. 1)
@@
These ideas are fleshed out further in more sophisticated stellar halo models (e.g. Bullock and Johnston, 2005; De Lucia and Helmi, 2008; Cooper et al., 2010; Monachesi et al., 2019; Fattahi et al., 2020), where the spatial distribution of accretion events can be explicitly tracked. Here, the dominance of a small number of massive progenitors is retained, but they are generally confined to the inner ($\lesssim 20-50$ kpc) halo, where they have been rapidly dragged inwards towards the centre of the host halo via dynamical friction. The small number of stellar halo building blocks and a wide range of merger configurations lead to a pronounced diversity of the resulting stellar halo density distributions (see e.g. Cooper et al., 2010), in contrast to the universality of the dark matter halo densities Navarro et al. (1996).
@@
Figure 1: Top Left: Galaxy luminosity function — the role of feedback (Silk and Mamon, 2012). Top Right: Schematic representation of a merger tree (Lacey and Cole, 1993). Time increases from top to bottom (present time $=t_{0}$, formation time $=t_{f}$), and the width of the branches represents the individual halo masses. Bottom panel: The mass-weighted distribution of destroyed dwarfs that contribute to the stellar haloes of Milky Way-mass systems. The thin coloured lines are for individual haloes, and the thick black line is the average. The thick dashed line shows the average profile for ‘quiescent’ haloes, which have not undergone a major merger since $z=2$. The dotted line shows the mass-weighted contribution of dark matter haloes, which have a much flatter distribution. This profile differs from the stellar mass owing to the steep decline in the stellar mass–halo mass relation at low halo masses ($\lesssim 10^{11}M_{\odot}$). Bottom insert: The (mass-weighted) number of dwarf progenitors that contribute to the total accreted stellar mass. Typically, $1-2$ destroyed dwarfs deposit the majority of accreted stellar mass onto the host halos [reproduced from Deason et al. 2016].
@@ 1.1.2 Pre-Gaia clues from observations
Number counts: A fundamental ‘zeroth order’ measurement of the Galactic stellar halo is to simply count the number of stars. As such, the characterization of the stellar halo density profile has been studied for several decades (e.g. Hartwick, 1987; Preston et al., 1991; Robin et al., 2000; Yanny et al., 2000; Jurić et al., 2008; Deason et al., 2011b; Xue et al., 2015). In recent years, however, it has been recognized that this is not simply a counting exercise, and the form of the stellar density profile in fact encodes key information pertaining to the assembly history of the halo (e.g. Bullock and Johnston, 2005; Cooper et al., 2010). Early work limited to the inner halo ($\lesssim 20$ kpc) described the stellar density as a simple power-law profile with radius ($\rho\propto r^{-\alpha}$, Robin et al. 2000, e.g.; Yanny et al. 2000, e.g.; Chen et al. 2001, e.g.; Siegel et al. 2002, e.g.; Jurić et al. 2008, e.g.). When the outer parts of the halo began to be mapped (largely thanks to wide-area photometric surveys such as SDSS), it was quickly recognised that this profile does not continue to larger radii. Indeed, Deason et al. (2011b) (also see Sesar et al. 2011) found that the stellar halo is better characterised by a ‘broken’ power-law with a significant steepening in the power-law (from $\alpha\sim 2.5$ to $\alpha\sim 4.5$) beyond $r\sim 25$ kpc (note that glimpses of the stellar halo break were seen earlier, see e.g. Saha, 1985; Watkins et al., 2009). In a follow-up paper, Deason et al. (2013a) provided a possible explanation for such a ‘broken’ profile (see Fig. 2). They used $N$-body simulations to argue that the debris from individual stellar halo progenitors often follow a broken power-law, where the break radius signifies the last apocenter of the orbit. In this scenario, a broken profile in the overall halo could imply two possible outcomes: (1) the stellar halo is dominated by one progenitor, or (2) several stellar halo progenitors all conspired to be accreted at the same time, with similar apocenters. The authors argued that the former scenario was more likely, owing to the contrived nature of the latter. However, the question remained, if the stellar halo was indeed dominated by one massive progenitor, how did we not know this already, and where is the additional observational evidence beyond the number counts?
@@
Figure 2: Left: Number counts of BHB stars from SDSS data. The blue line is the best-fit broken power-law, and the red line is the best-fit single power-law (Deason et al., 2011b). Right: Example of an individual accretion event from the Bullock and Johnston (2005) simulation suite. The panels show the radial velocity (top) and density profile (bottom) as a function of radius. The dashed line indicates the ‘break-radius’ of the best-fit broken power law that coincides with the apocenter of the debris. The black box indicates the ‘local’ region of the debris, which has a high density and significant radial velocity $v_{r}$ [reproduced from Deason et al. 2013a].
@@
In the influential work by Venn et al. (2004) the authors compared chemical abundances of halo stars to surviving dwarf spheroidal satellite galaxies. They found that their chemical signatures are remarkably different, perhaps ruling out the common lore that the stellar halo is built up from the debris of many low-mass galaxies! However, importantly, the authors note that mergers with higher mass dwarf galaxies cannot be ruled out, as the [$\alpha$/Fe] vs [Fe/H] plane for the halo stars exhibits a relatively metal-rich knee (although less metal-rich than the disc stars). An additional conundrum is that the chemistry of the halo stars also does not resemble the Large Magellanic Cloud (LMC) or Sagittarius (Sgr), which are the two of the three most massive dwarf galaxies ($M_{\rm star}\sim 10^{8-9}M_{\odot}$) in the Milky Way. The $\Lambda$CDM simulation community was swift to clear this up: the stellar halo models predict that the destroyed dwarf progenitors were accreted early, and thus have star formation histories different to the late-time Sgr and LMC type objects that can continue to form stars (Font et al., 2006). Font et al. (2006) showed explicitly how the $\Lambda$CDM accretion plus chemical evolution models can match the Milky Way data. In summary, from the chemical perspective, the dominance of early, massive dwarf galaxy progenitors seems to fit the bill.
A final intriguing discovery in the chemical analysis of halo stars came from the work by Nissen and Schuster (2010). These authors used precise abundance ratios to show that there are two distinct halo populations in the solar neighbourhood. One ‘high-$\alpha$’ sequence can perhaps be attributed to an ‘in-situ’ halo component (see Section 4.1), while the other ‘low-$\alpha$’ sequence resembles the classical ‘accreted’ component. However, what was remarkable about this analysis was just how narrow and ordered the chemical sequence of the accreted stars was in the [$\alpha$/Fe] vs. [Fe/H] plane. Presumably, a hodge-podge of several different progenitors with different masses and SFH would produce a messier outcome. Alternatively, at least in the solar neighborhood, there really is one progenitor that dominates the accreted halo.
@@
In the pre-Gaia era, accurate space velocity measurements for distant objects were hard to come by. As a result, outside of the Solar neighborhood, the view of the halo’s orbital properties remained rather blurred. For Galactic GCs it is possible to beat down the proper motion errors by averaging over large numbers of member stars. Thus, through early heroic efforts first estimates of the Milky Way’s GC orbital properties were obtained (see e.g. Gnedin and Ostriker, 1997; Dinescu et al., 1999). These studies found the GC velocity ellipsoid to be broadly consistent with that measured using stellar kinematics in the Solar neighborhood. In addition, Dinescu et al. (1999) found systematic — albeit sometimes rather subtle — differences in the orbital properties of the GCs as a function of their age, chemistry, and location in the Galaxy. For example, in their sample, the clusters’ orbital eccentricity exhibited dependence on metallicity, with the metal-poorer GCs reaching higher eccentricities and the metal-richer ones tending towards more circular orbits.
@@
A very different orbital structure forms in the stellar halo in the series of experiments discussed in the visionary work by Amorisco (2017). In their toy $N$-body models, massive accreted satellites lose most of their angular momentum in a rapid tidal disruption, driven in part by strong dynamical friction. Amorisco (2017) demonstrate (see their Figure 11) that the resulting debris has a strong radial anisotropy, reaching close to $\beta\approx 1$ at the approximate location of the Sun (see also right-panel of Fig. 2). This orbital radialization reported by Amorisco (2017) for massive satellites went contrary to the previously established intuition where dynamical friction helped to circularise the satellite’s orbit instead (see e.g. Jiang and Binney, 2000; Nipoti, 2017). While hints of orbital radialization had been seen already by Barnes (1988), the mechanics of the angular momentum loss during the accretion of massive satellites were finally explained in Vasiliev et al. (2022). Thus, this relatively new explanation of massive satellite radialization can naturally account for the high $\beta$ values found in the solar neighbourhood if the inner halo is dominated by the debris from massive systems.
@@
Finally, several trailblazing studies managed to demonstrate the principles of Galactic Archaeology in action, even with the scarce data available before Gaia. These early pioneers combine chemistry and orbital information to detect individual accretion events in the Solar neighborhood. Helmi et al. (1999) discover a stellar stream from a small dwarf galaxy passing near the Sun, in a clear example of what is to become a routine in the Gaia era. Equally striking are the first likely hints of the tidal debris from the last significant merger in the works of Dinescu (2002) and Brook et al. (2003). Both use the chemo-kinematic dataset by Chiba and Beers (2000). The study of Dinescu (2002) finds evidence for a retrograde disruption event, while Brook et al. (2003) interprets the observations as a signal from a recent accretion of a dwarf on a polar orbit.

### v2

@@ Abstract
The Gaia mission has revolutionized our view of the Milky Way and its satellite citizens. The field of Galactic Archaeology has been piecing together the formation and evolution of the Galaxy for decades, and we have made great strides, with often limited data, towards discovering and characterizing the subcomponents of the Galaxy and its building blocks. Now, the exquisite 6D phase-space plus chemical information from Gaia and its complementary spectroscopic surveys has handed us a plethora of data to pore over as we move towards a quantitative rather than qualitative view of the Galaxy and its progenitors. We review the state of the field in the post-Gaia era, and examine the key lessons that will dictate the future direction of Galactic halo research.
@@ 1 Introduction
Gaia is an astrometric mission with the lofty goal of producing an all-sky phase-space map of the Galaxy (Gaia Collaboration et al., 2016b). Positions on the sky, parallax, and proper motions are measured for all stars down to a limiting magnitude of $G\approx 21$. In addition, a (bright) subset of stars has measured radial velocities and stellar parameters (such as effective temperature and surface gravity, see Recio-Blanco et al. 2023 for details of the parametrization of the Gaia RVS data). As of June 2024, we are now on the third Gaia data release with the fourth release likely arriving in 2025/26. Each subsequent data release provides data over a larger time baseline, and hence more precise astrometry, as well as additional data products, such as variability information and source classifications. To compensate for the bright limit for accurate Gaia spectroscopic measurements, several wide-field spectroscopic surveys have (or will) complemented the Gaia data to provide radial velocities and chemical measurements, e.g. SDSS/SEGUE (Yanny et al., 2009), LAMOST (Cui et al., 2012), APOGEE (Majewski et al., 2017), GALAH (Buder et al., 2021), H3 (Conroy et al., 2019), DESI (Cooper et al., 2023), WEAVE (Jin et al., 2023), 4MOST (de Jong et al., 2019). These auxiliary datasets are particularly important for the halo stars, which traverse to large distances and are thus often relatively faint. The era of Gaia is truly data-heavy! The fruits of this labour will last for many years to come, and we are only at the beginning stages of our digestion and understanding of what this detailed mapping of the Milky Way contains.
The purpose of this review is to provide the current state of understanding of the Galactic halo and the revelations and surprises from the Gaia data boon. This review does not intend to provide a complete history and consensus, as the number of papers in this field has exploded in recent years. Instead, we focus on the main findings, in our humble opinion, and look ahead to the future of this field over the coming decade. There are several articles in this special issue focused on “Gaia discoveries”. As such, the in-depth exploration of stellar streams, globular clusters and Galactic dynamics with Gaia is covered elsewhere and we refer the interested reader to related reviews on those topics.
@@ 1.1 A dominant stellar halo progenitor?
One of the main aims of the Gaia mission is to piece together the building blocks of the Galactic halo. This ambitious goal was perhaps mainly expected to reveal a plethora of streams, clouds, and debris from a multitude of progenitors (see e.g. Helmi and de Zeeuw, 2000). However, one of the clearest and most decisive discoveries from the early Gaia data releases was that in fact the (inner, $r\lesssim 20$ kpc) stellar halo is dominated by one dwarf galaxy progenitor (see Section 2). As we will discuss below, in hindsight this was not a surprising discovery!
@@ 1.1.1 Predictions from $\Lambda$CDM
The fundamental mass function of dark matter haloes ($dn/dM$ — i.e. how many haloes there are of different mass) follows a power-law profile $dn/dM\propto M^{-1.9}$ (e.g. Moore et al., 1999; Jenkins et al., 2001; Gao et al., 2004; Tinker et al., 2008). Thus, the general expectation is that there are many more smaller lumps than bigger lumps — a prediction that lies at the heart of the hierarchical $\Lambda$CDM paradigm. A consequence of the mass function is that haloes like our own Milky Way engulf many more smaller mass objects than more massive ones over time. For example, cosmological $N$-body simulations (Fakhouri et al., 2010) predict that a Milky Way-mass halo ($10^{12}M_{\odot}$) typically undergoes $N\sim 30$ minor mergers (mass-ratio between 1:3 and 1:100) and $N\sim 3$ major mergers (mass-ratio $>1:3$) since redshift $z=10$.
@@
When galaxies enter the equation, things become more complicated as the efficiency of star formation is not equal amongst haloes of different mass. The inefficiency of massive galaxies in the centres of clusters (mainly due to AGN feedback, e.g. Croton et al. 2006; Hopkins et al. 2006; Fabian 2012) and dwarf galaxies (supernova feedback plus reionization, e.g. Dekel and Silk 1986; Efstathiou 1992; Bullock et al. 2000; Pontzen and Governato 2012) at forming stars means that galaxies like the Milky Way lead the pack — we live in the most efficient star-making galaxy environment (see e.g. Figure 2 of Wechsler and Tinker, 2018). The relation between stellar mass and dark matter mass (aptly known as the ‘stellar mass-halo mass’ or SMHM relation), has important consequences for the growth of stellar haloes through the accretion of lower-mass objects. The SMHM has a very steep decline at low halo masses ($\lesssim 10^{11}M_{\odot}$, see e.g. Figure 6 of Bullock and Boylan-Kolchin 2017). Thus, although there may be hundreds of dark matter subhaloes accreted by the galaxy, if these have very few stars (or no stars at all!) then they have an insignificant contribution to the accreted stellar halo. This idea was recognized in Purcell et al. (2007), who used analytic prescriptions for dark matter halo growth and galaxy formation models, and later by Deason et al. (2016) using $N$-body simulations and analytic SMHM relations. In short, the contribution of stellar mass to the galactic stellar halo is dominated by a small number (1-2) of relatively massive dwarfs ($M_{\rm DM}\sim 10^{11}M_{\odot}$, $M_{\rm star}\sim 10^{8-9}M_{\odot}$). The contribution of lower mass dwarfs, while potentially important for dark matter growth, is relatively insignificant for the total stellar halo mass (see Fig. 1)
@@
These ideas are fleshed out further in more sophisticated stellar halo models (e.g. Bullock and Johnston, 2005; De Lucia and Helmi, 2008; Cooper et al., 2010; Monachesi et al., 2019; Fattahi et al., 2020; Horta et al., 2023a), where the spatial distribution of accretion events can be explicitly tracked. Here, the dominance of a small number of massive progenitors is retained, but they are generally confined to the inner ($\lesssim 20-50$ kpc) halo, where they have been rapidly dragged inwards towards the centre of the host halo via dynamical friction. The small number of stellar halo building blocks and a wide range of merger configurations lead to a pronounced diversity of the resulting stellar halo density distributions (see e.g. Cooper et al., 2010), in contrast to the universality of the dark matter halo densities Navarro et al. (1996).
@@
Figure 1: Top Left: Galaxy luminosity function — the role of feedback (Silk and Mamon, 2012). Top Right: Schematic representation of a merger tree (Lacey and Cole, 1993). Time increases from top to bottom (present time $=t_{0}$, formation time $=t_{f}$), and the width of the branches represents the individual halo masses. Bottom panel: The mass-weighted distribution of destroyed dwarfs that contribute to the stellar haloes of Milky Way-mass systems (Deason et al., 2016). The thin coloured lines are for individual haloes, and the thick black line is the average. The thick dashed line shows the average profile for ‘quiescent’ haloes, which have not undergone a major merger since $z=2$. The dotted line shows the mass-weighted contribution of dark matter haloes, which have a much flatter distribution. This profile differs from the stellar mass owing to the steep decline in the stellar mass–halo mass relation at low halo masses ($\lesssim 10^{11}M_{\odot}$). Bottom insert: The (mass-weighted) number of dwarf progenitors that contribute to the total accreted stellar mass. Typically, $1-2$ destroyed dwarfs deposit the majority of accreted stellar mass onto the host halos [reproduced from Silk and Mamon 2012; Lacey and Cole 1993; Deason et al. 2016].
@@ 1.1.2 Pre-Gaia clues from observations
Number counts: A fundamental ‘zeroth order’ measurement of the Galactic stellar halo is to simply count the number of stars. As such, the characterization of the stellar halo density profile has been studied for several decades (e.g. Hartwick, 1987; Preston et al., 1991; Robin et al., 2000; Yanny et al., 2000; Jurić et al., 2008; Deason et al., 2011b; Xue et al., 2015). In recent years, however, it has been recognized that this is not simply a counting exercise, and the form of the stellar density profile in fact encodes key information pertaining to the assembly history of the halo (e.g. Bullock and Johnston, 2005; Cooper et al., 2010). Early work limited to the inner halo ($\lesssim 20$ kpc) described the stellar density as a simple power-law profile with radius ($\rho\propto r^{-\alpha}$, Robin et al. 2000, e.g.; Yanny et al. 2000, e.g.; Chen et al. 2001, e.g.; Siegel et al. 2002, e.g.; Jurić et al. 2008, e.g.). When the outer parts of the halo began to be mapped (largely thanks to wide-area photometric surveys such as SDSS), it was quickly recognized that this profile does not continue to larger radii. Indeed, Deason et al. (2011b) (also see Sesar et al. 2011) found that the stellar halo is better characterised by a ‘broken’ power-law with a significant steepening in the power-law (from $\alpha\sim 2.5$ to $\alpha\sim 4.5$) beyond $r\sim 25$ kpc (note that glimpses of the stellar halo break were seen earlier, see e.g. Saha, 1985; Watkins et al., 2009). In a follow-up paper, Deason et al. (2013a) provided a possible explanation for such a ‘broken’ profile (see Fig. 2). They used $N$-body simulations to argue that the debris from individual stellar halo progenitors often follow a broken power-law, where the break radius signifies the last apocentre of the orbit. In this scenario, a broken profile in the overall halo could imply two possible outcomes: (1) the stellar halo is dominated by one progenitor, or (2) several stellar halo progenitors all conspired to be accreted at the same time, with similar apocentres. The authors argued that the former scenario was more likely, owing to the contrived nature of the latter. However, the question remained, if the stellar halo was indeed dominated by one massive progenitor, how did we not know this already, and where is the additional observational evidence beyond the number counts?
@@
Figure 2: Left: Number counts of BHB stars from SDSS data. The blue line is the best-fit broken power-law, and the red line is the best-fit single power-law (Deason et al., 2011b). Right: Example of an individual accretion event from the Bullock and Johnston (2005) simulation suite. The panels show the radial velocity (top) and density profile (bottom) as a function of radius. The dashed line indicates the ‘break-radius’ of the best-fit broken power law that coincides with the apocentre of the debris. The black box indicates the ‘local’ region of the debris, which has a high density and significant radial velocity $v_{r}$ [reproduced from Deason et al. 2013a].
@@
In the influential work by Venn et al. (2004) the authors compared chemical abundances of halo stars to surviving dwarf spheroidal satellite galaxies. They found that their chemical signatures are remarkably different, perhaps ruling out the common lore that the stellar halo is built up from the debris of many low-mass galaxies! However, importantly, the authors note that mergers with higher mass dwarf galaxies cannot be ruled out, as the [$\alpha$/Fe] vs [Fe/H] plane for the halo stars exhibits a relatively metal-rich knee (although less metal-rich than the disc stars). An additional conundrum is that the chemistry of the halo stars also does not resemble the Large Magellanic Cloud (LMC) or Sagittarius (Sgr), which are two of the three most massive dwarf galaxies ($M_{\rm star}\sim 10^{8-9}M_{\odot}$) in the Milky Way. The $\Lambda$CDM simulation community was swift to clear this up: the stellar halo models predict that the destroyed dwarf progenitors were accreted early, and thus have star formation histories different to the late-time Sgr and LMC type objects that can continue to form stars (Font et al., 2006). Font et al. (2006) showed explicitly how the $\Lambda$CDM accretion plus chemical evolution models can match the Milky Way data. In summary, from the chemical perspective, the dominance of early, massive dwarf galaxy progenitors seems to fit the bill.
A final intriguing discovery in the chemical analysis of halo stars came from the work by Nissen and Schuster (2010). These authors used precise abundance ratios to show that there are two distinct halo populations in the solar neighbourhood. One ‘high-$\alpha$’ sequence can perhaps be attributed to an ‘in-situ’ halo component (see Section 4.1), while the other ‘low-$\alpha$’ sequence resembles the classical ‘accreted’ component. However, what was remarkable about this analysis was just how narrow and ordered the chemical sequence of the accreted stars was in the [$\alpha$/Fe] vs. [Fe/H] plane. Presumably, a hodge-podge of several different progenitors with different masses and SFH would produce a messier outcome. Alternatively, at least in the solar neighourhood, there really is one progenitor that dominates the accreted halo.
@@
In the pre-Gaia era, accurate space velocity measurements for distant objects were hard to come by. As a result, outside of the Solar neighourhood, the view of the halo’s orbital properties remained rather blurred. For Galactic GCs it is possible to beat down the proper motion errors by averaging over large numbers of member stars. Thus, through early heroic efforts first estimates of the Milky Way’s GC orbital properties were obtained (see e.g. Gnedin and Ostriker, 1997; Dinescu et al., 1999). These studies found the GC velocity ellipsoid to be broadly consistent with that measured using stellar kinematics in the Solar neighourhood. In addition, Dinescu et al. (1999) found systematic — albeit sometimes rather subtle — differences in the orbital properties of the GCs as a function of their age, chemistry, and location in the Galaxy. For example, in their sample, the clusters’ orbital eccentricity exhibited dependence on metallicity, with the metal-poorer GCs reaching higher eccentricities and the metal-richer ones tending towards more circular orbits.
@@
A very different orbital structure forms in the stellar halo in the series of experiments discussed in the visionary work by Amorisco (2017). In their toy $N$-body models, massive accreted satellites lose most of their angular momentum in a rapid tidal disruption, driven in part by strong dynamical friction. Amorisco (2017) demonstrate (see their Figure 11) that the resulting debris has a strong radial anisotropy, reaching close to $\beta\approx 1$ at the approximate location of the Sun (see also right-panel of Fig. 2). This orbital radialization reported by Amorisco (2017) for massive satellites went contrary to the previously established intuition where dynamical friction helped to circularize the satellite’s orbit instead (see e.g. Jiang and Binney, 2000; Nipoti, 2017). While hints of orbital radialization had been seen already by Barnes (1988), the mechanics of the angular momentum loss during the accretion of massive satellites were finally explained in Vasiliev et al. (2022). Thus, this relatively new explanation of massive satellite radialization can naturally account for the high $\beta$ values found in the solar neighbourhood if the inner halo is dominated by the debris from massive systems.
@@
Finally, several trailblazing studies managed to demonstrate the principles of Galactic Archaeology in action, even with the scarce data available before Gaia. These early pioneers combine chemistry and orbital information to detect individual accretion events in the Solar neighourhood. Helmi et al. (1999) discover a stellar stream from a small dwarf galaxy passing near the Sun, in a clear example of what is to become a routine in the Gaia era. Equally striking are the first likely hints of the tidal debris from the last significant merger in the works of Dinescu (2002) and Brook et al. (2003). Both use the chemo-kinematic dataset by Chiba and Beers (2000). The study of Dinescu (2002) finds evidence for a retrograde disruption event, while Brook et al. (2003) interprets the observations as a signal from a recent accretion of a dwarf on a polar orbit.